General Data: 1. Sales are 25% cash, 75% on credit. 2. Of the credit sales, 80% are collected in the month following the month of sale and 20% in the second month following the sale. 3, Gross Profit margin on sales averages 25% i.e., the COGS is 75% of sales. 4. All inventory purchases are paid during the month in which they are made 5. The store follows the policy of purchasing enough inventory each month to cover the following months sales at cost 6. A minimum cash balance of $5,000 is to be maintained by the store 7. A $15,000 dividend payment will be made in January 8. The cash balance as of July 31 is $10,400 9 Sales Budget August $14,000 November $38,000 September 26,000 December 48,000 October 22,000 January 18,000 February 16,000 10. Monthly Expenses
media%2Fa9f%2Fa9f167e0-ba47-4071-851c-b8
media%2F579%2F579a3ce7-ee0b-40c0-99a8-60
media%2Fa8c%2Fa8c9a7a4-ac62-4139-93c7-94
media%2F80f%2F80f6604d-241d-4bc5-9ccc-7b
media%2F634%2F634a16a5-7499-4fb8-b7a3-88
media%2Face%2Facea8b3c-7d05-472e-9e0f-ee
media%2F596%2F596bfc95-51a2-4e84-b19f-a6
Answer #1

Answer to Question 1.

Total Cash Inflow in January = (25% of January Sales) + (80% of December Credit Sales) + (20% of November Credit Sales)

December Credit Sales = $48,000 * 75% = $36,000
November Credit Sales = $38,000 * 75% = $28,500

Total Cash Inflow in January = (25% of $18,000) + (80% of $36,000) + (20% of $28,500)
Total Cash Inflow in January = $4,500 + $28,800 + $5,700
Total Cash Inflow in January = $39,000

Answer to Question 2.

Total Expected Cash Outflow for Inventory in December = Cost of Inventory purchased in December
Cost of Inventory purchased in December = Cost of Inventory to be sold in January
Cost of Inventory purchased in December = (75% of $18,000)
Cost of Inventory purchased in December = $13,500

Total Expected Cash Outflow for Inventory in December = $13,500

Answer to Question 3.

Total Expected cash Outflow for January = Expected Cash Outflow for Inventory + Dividend Payment + Wages Payment + Rent Payment +Other Monthly Expense

Total Expected Cash Outflow for Inventory in January = Cost of Inventory purchased in January
Cost of Inventory purchased in January = Cost of Inventory to be sold in February
Cost of Inventory purchased in January = (75% of $16,000)
Cost of Inventory purchased in January = $12,000

Total Expected Cash Outflow for Inventory in January = $12,000

Total Expected cash Outflow for January = $12,000 + $15,000 + $1,400 + $400 + (1% of $18,000)
Total Expected cash Outflow for January = $12,000 + $15,000 + $1,400 + $400 + $180
Total Expected cash Outflow for January = $28,980

Answer to Question 4.

Accounts Receivable, January 31 = January Credit Sales + (20% of December Credit Sales)

January Credit Sales = $18,000 * 75% = $13,500
December Credit Sales = $48,000 * 75% = $36,000

Accounts Receivable, January 31 = $13,500 + (20% of $36,000)
Accounts Receivable, January 31 = $13,500 + $7,200
Accounts Receivable, January 31 = $20,700